# Generalization Error of the Tilted Empirical Risk

### v2

@@ 7 Related Works
Generalization Error Analysis: Different approaches have been applied to study the generalization error of general learning problems under empirical risk minimization, including VC dimension-based, Rademacher complexity, PAC-Bayesian, stability and information-theoretic bounds. In this section, we discuss the related works about Uniform and information-theoretic bounds. More related works for generalization error analysis are discussed in Appendix B.
Uniform Bounds: Uniform bounds (or VC bounds) are proposed by  Vapnik and Chervonenkis (1971); Bartlett et al. (1998, 2019). For any class of functions $\mathcal{F}$ of VC dimension $d$, with probability at least $1-\delta$ the generalization error is $O\big{(}(d+\log(1/\delta))^{1/2}n^{-1/2}\big{)}$. This bound depends solely on the VC dimension of the function class and on the sample size; in particular, it is independent of the learning algorithm.
Information-theoretic bounds: Russo and Zou (2019); Xu and Raginsky (2017) propose using the mutual information between the input training set and the output hypothesis to upper bound the expected generalization error. Multiple approaches have been proposed to tighten the mutual information-based bound: Bu et al. (2020) provide tighter bounds by considering the individual sample mutual information; Asadi et al. (2018); Asadi and Abbe (2020) propose using chaining mutual information; and Steinke and Zakynthinou (2020); Hafez-Kolahi et al. (2020); Aminian et al. (2020, 2021b) provide different upper bounds on the expected generalization error based on the linear empirical risk framework.
The aforementioned approaches are applied to study the generalization error in the linear empirical risk framework. To our knowledge, the generalization error of the tilted empirical risk minimization from information theoretical and uniform approach perspectives has not been explored.
8 Conclusion and Future Work
@@
In this paper, we study the tilted empirical risk minimization, as proposed by Li et al. (2021). In particular, we established an upper and lower bound on the tilted generalization error of the tilted empirical risk through uniform and information-theoretic approaches, obtaining theoretical guarantees that the convergence rate is $O(1/\sqrt{n})$ under bounded loss functions for negative and positive tilts. Furthermore, we provide an upper bound on the titled generalization error for the unbounded loss function. We also study the tilted generalization error under distribution shift in the training dataset due to noise or outliers, where we discussed the generalization and robustness trade-off. Additionally, we explore the KL-regularized tilted empirical risk minimization, where the solution involves the tilted Gibbs posterior, and we derive a parametric upper bound on this minimization with a convergence rate of $O(1/n)$ under some conditions.
Our current results are applicable to scenarios with bounded loss functions and in scenarios with negative tilting ($\gamma<0$) when the possibly unbounded loss function has a bounded second moment. However, our current results for the asymptotic regime $\gamma\rightarrow\infty$, where the tilted empirical risk is equal to maximum loss, are vacuous. Therefore, studying the generalization performance of tilted empirical risk minimization under unbounded loss functions for positive tilting, and obtaining informative bounds in the asymptotic regime $\gamma\rightarrow\infty$, are planned as future work. Moreover, we intend to apply other approaches from the literature, such as those discussed by Aminian et al. (2023), to derive further bounds on the tilted generalization error in the mean-field regime.
Acknowledgements
@@
Gholamali Aminian, Gesine Reinert and Samuel N. Cohen acknowledge the support of the UKRI Prosperity Partnership Scheme (FAIR) under EPSRC Grant EP/V056883/1 and the Alan Turing Institute. Gesine Reinert is also supported in part by EPSRC grants EP/W037211/1 and EP/R018472/1. Samuel N. Cohen also acknowledges the support of the Oxford–Man Institute for Quantitative Finance. Amir R. Asadi is supported by Leverhulme Trust grant ECF-2023-189 and Isaac Newton Trust grant 23.08(b).
Appendix A Overview of Main Results
An overview of our main results is provided in Fig. 1. All notations are summarized in Table 1.
@@ Appendix A Overview of Main Results
$\overline{\mathrm{R}}_{\gamma}(h,P_{S})$ Tilted population risk $\mathfrak{E}_{\gamma}(\mu)$ Excess risk
Tilted Generalization Error Bounded Loss Function ($\gamma\in\mathbb{R}$) Section 3 Information-theoretic Bound Upper bound (Theorem 3.13) & Lower bound (Theorem 3.14) Uniform Bound Upper bound (Theorem 3.3) & Lower bound (Theorem 3.4) Unbounded Loss Function ($\gamma\in\mathbb{R}_{<0}$) Section 4 Information-theoretic Bound Upper bound (Theorem E.4) & Lower bound (Theorem E.6) Uniform Bound Upper bound (Theorem 4.2) & Lower bound (Theorem 4.3) Robustness of TERM ($\gamma\in\mathbb{R}_{<0}$) Section 5 Uniform Bound Upper bound (Proposition 5.3) & Lower bound (Theorem 5.4) KL-Regularized TERM ($\gamma\in\mathbb{R}$) Section 6 Information-theoretic Bound Upper bound (Theorem 6.3)
Figure 1: Overview of the main results
@@ Appendix B Other Related Works
This section details related works about Tilted empirical risk minimization, the Rademacher complexity, stability and PAC-Bayesian bounds, as well as related work on unbounded loss functions.
@@
Rademacher Complexity Bounds: This approach is a data-dependent method to provide an upper bound on the generalization error based on the Rademacher complexity of the function class $\mathcal{H}$, see (Bartlett and Mendelson, 2002; Golowich et al., 2018). Bounding the Rademacher complexity involves the model parameters. Typically, in Rademacher complexity analysis, a symmetrization technique is used which can be applied to the empirical risk, but not directly to the TER.
@@
Stability Bounds: Stability-based bounds for generalization error are given in Bousquet and Elisseeff (2002a); Bousquet et al. (2020); Mou et al. (2017); Chen et al. (2018); Aminian et al. (2023). For stability analysis, the key tool is Lemma 7 in Bousquet and Elisseeff (2002b), which is based on ERM linearity. Therefore, we can not apply stability analysis to TER directly.
@@
PAC-Bayes bounds: First proposed by Shawe-Taylor and Williamson (1997); McAllester (1999) and McAllester (2003), PAC-Bayesian analysis provides high probability bounds on the generalization error in terms of the KL divergence between the data-dependent posterior induced by the learning algorithm and a data-free prior that can be chosen arbitrarily (Alquier, 2021). There are multiple ways to generalize the standard PAC-Bayesian bounds, including using information measures other than KL divergence (Alquier and Guedj, 2018; Bégin et al., 2016; Hellström and Durisi, 2020; Aminian et al., 2021b) and considering data-dependent priors (Rivasplata et al., 2020; Catoni, 2007; Dziugaite and Roy, 2018; Ambroladze et al., 2007). However, this method has not been applied to TER to provide generalization error bounds.
Tilted Empirical Risk Minimization: The TERM algorithm for machine learning is proposed by Li et al. (2021), and good performance of the TERM under outlier and noisy label scenarios for negative tilting ($\gamma<0$) and under imbalance and fairness constraints for positive tilting ($\gamma>0$) is demonstrated. Inspired by TERM, Wang et al. (2023) propose a class of new tilted sparse additive models based on the Tikhonov regularization scheme. Their results have some limitations. First, in (Wang et al., 2023, Theorem 3.3) the authors derive an upper bound for $\lambda=n^{-\zeta}$ where $\zeta<-1/2$ and $\lambda$ are the regularization parameters in (Wang et al., 2023, Eq.4). This implies $\lambda\rightarrow\infty$ as $n\rightarrow+\infty$, which is impractical. Finally, the analysis in (Wang et al., 2023) assumes that both the loss function and its derivative are bounded. Therefore, it can not be applied to the unbounded loss function scenario. Furthermore, we consider KL regularization, which is different from the Tikhonov regularization scheme with the sparsity-induced $\ell_{1,2}$-norm regularizer as introduced in (Wang et al., 2023). Therefore, our current results do not cover the learning algorithm in [9]. Zhang et al. (2023) studied the TERM as a target function to improve the robustness of estimators. The application of TERM in federated learning is also studied, in (Li et al., 2023b; Zhang et al., 2022). The authors in (Lee et al., 2020), propose an upper bound on the Entropic risk function generalization error via the representation of coherent risk function and using the Rademacher complexity approach. However, their approach is limited to negative tilt and bounded loss function. Although rich experiments are given by Li et al. (2021) for the TERM algorithm in different applications, the generalization error of the TERM has not yet been addressed for unbounded loss functions.
Unbounded loss functions: Some works studied the generalization error under unbounded loss functions via the PAC-Bayesian approach. Losses with heavier tails are studied by Alquier and Guedj (2018) where probability bounds (non-high probability) are developed. Using a different estimator than empirical risk, PAC-Bayes bounds for losses with bounded second and third moments are developed by Holland (2019). Notably, their bounds include a term that can increase with the number of samples $n$. Kuzborskij and Szepesvári (2019) and Haddouche and Guedj (2022) also provide bounds for losses with a bounded second moment. The bounds in Haddouche and Guedj (2022) rely on a parameter that must be selected before the training data is drawn. Information-theoretic bounds based obased on the second moment of $\sup_{h\in\mathcal{H}}|\ell(h,Z)-\mathbb{E}[\ell(h,\tilde{Z})]|$ are derived in Lugosi and Neu (2022, 2023). In contrast, our second moment assumption is more relaxed, being based on the expected version with respect to the distribution over the hypothesis set and the data-generating distribution. In our paper, we provided a generalization error bound on the tilted empirical risk via uniform and information-theoretic approaches under a bounded second-moment assumption.
@@ Appendix C Technical Tools
We first define the functional linear derivative as in Cardaliaguet et al. (2019).
@@
Suppose that $X>0$ and $\gamma<0$, then we have
$\mathrm{Var}(\exp(\gamma X))\leq\gamma^{2}\mathrm{Var}(X).$ (32)
@@
By the mean value theorem, for each realisation $X(\omega)$ of $X$ for an element $\omega$ of its underlying probability space there is a value $c(\omega)$ in the interval between $X(\omega)$ and $\mathbb{E}[X]$ such that
$\exp(\gamma X(\omega))-\exp(\gamma\mathbb{E}[X])=\gamma(X-\mathbb{E}[X])\exp(\gamma c(\omega)).$
As $X>0$ we have $c(\omega)>0$. Moreover,
$\begin{split}\mathrm{Var}(\exp(\gamma X))&=\mathbb{E}[(\exp(\gamma X)-\mathbb{E}[\exp(\gamma X)])^{2}]\\ &\overset{(a)}{\leq}\mathbb{E}[(\exp(\gamma X)-\exp(\gamma\mathbb{E}[X]))^{2}]\\ &\overset{(b)}{=}\mathbb{E}[\gamma^{2}\exp(2\gamma c)(X-\mathbb{E}[X])^{2}]\\ &\overset{(c)}{\leq}\gamma^{2}\mathrm{Var}(X),\end{split}$
where $(a)$, $(b)$ and $(c)$ follow from the minimum mean square representation, the mean-value theorem and the negativity of $\gamma X$, respectively. ∎
@@
Lemma C.5 (McDiarmid’s inequality (McDiarmid, 1989)).
Let $F:\mathcal{Z}^{n}\mapsto\mathbb{R}$ be any measurable function for which there exists constants $c_{i},i=1,\cdots,m)$ such that
$\sup_{S\in\mathcal{Z}^{n},\tilde{Z}_{i}\in\mathcal{Z}}|F(S)-F(S_{(i)}|\leq c_{i},$
where $S_{(i)}=\{Z_{1},\cdots,\tilde{Z}_{i},\cdots,Z_{n}\},$ is the replace-one sample dataset and $\tilde{Z}_{i}$ is an i.i.d. sample with respect to all $Z_{i}$ for $i\in[n]$. Then the following inequality holds with probability at least $(1-\delta)$ under $P_{S}$,
$|F(S)-\mathbb{E}_{P_{S}}[F(S)]|\leq\sqrt{\big{(}\sum_{i=1}^{n}c_{i}^{2}\big{)}\frac{\log(1/\delta)}{2}}.$
@@
For a positive random variable, $Z>0$, suppose $\mathbb{E}[Z^{2}]\leq\eta$ Then, the following inequality holds,
$\mathbb{E}[Z]\leq\eta^{1/2}.$
@@
Suppose $\mathbb{E}[X^{2}]<\infty$. Then, for $X>0$ and $\gamma<0$ the following inequality holds,
$0\leq\mathbb{E}[X]-\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}\leq\frac{-\gamma}{2}\mathbb{E}[X^{2}].$
@@
The left inequality follows from Jensen’s inequality applied to $f(x)=\log{(x)}$. For the right inequality, we have for $\gamma X<0$,
$e^{\gamma X}\leq 1+\gamma X+\frac{1}{2}(\gamma X)^{2}.$ (35)
Therefore, we have
$\begin{split}\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}&\geq\frac{1}{\gamma}\log{\mathbb{E}\left[1+\gamma X+\frac{1}{2}\gamma^{2}X^{2}\right]}\\ &=\frac{1}{\gamma}\log{\left(1+\gamma\mathbb{E}[X]+\frac{1}{2}\gamma^{2}\mathbb{E}[X^{2}]\right)}\\ &\geq\frac{1}{\gamma}\left(\gamma\mathbb{E}[X]+\frac{1}{2}\gamma^{2}\mathbb{E}[X^{2}]\right)\\ &=\mathbb{E}[X]+\frac{\gamma}{2}\mathbb{E}[X^{2}].\end{split}$
@@
Lemma C.12 (Talagrand’s contraction lemma (Shalev-Shwartz and Ben-David, 2014)).
@@
Lemma C.14 (Lemma 1 in (Xu and Raginsky, 2017)).
@@
Lemma C.16 (Kantorovich-Rubenstein duality of total variation distance, see  (Polyanskiy and Wu, 2022)).

### v3

@@ 7 Related Works
In this section, we discuss related works on tilted empirical risk minimization and generalization error under an unbounded loss function. More related works for generalization error analysis are discussed in Appendix B.
Tilted Empirical Risk Minimization: The TERM algorithm for machine learning is proposed by Li et al. (2021), where good performance of the TERM under outlier and noisy label scenarios for negative tilting ($\gamma<0$), and, under imbalance and fairness constraints, for positive tilting ($\gamma>0$), is demonstrated. Zhang et al. (2023) study the TERM as a target function to improve the robustness of estimators. The application of TERM in federated learning is also studied, in Li et al. (2023b); Zhang et al. (2022).
Inspired by TERM, Wang et al. (2023) propose a class of new tilted sparse additive models based on the Tikhonov regularization scheme. Their results have some limitations. First, in (Wang et al., 2023, Theorem 3.3) the authors derive an upper bound for $\lambda=n^{-\zeta}$ where $\zeta<-1/2$ and $\lambda$ are the regularization parameters in (Wang et al., 2023, Eq.4). This implies $\lambda\rightarrow\infty$ as $n\rightarrow+\infty$, which is impractical. As the analysis in (Wang et al., 2023) assumes that both the loss function and its derivative are bounded, it can not be applied to the unbounded loss function scenario. Furthermore, we consider KL regularization, which is different from the Tikhonov regularization scheme with the sparsity-induced $\ell_{1,2}$-norm regularizer as introduced in Wang et al. (2023). Therefore, our results do not cover the learning algorithm in Wang et al. (2023). Lee et al. (2020) propose an upper bound on the entropic risk function generalization error via a representation of a coherent risk function and using the Rademacher complexity approach. However, their approach is limited to bounded loss functions.
Tilted risk for off-policy learning and evaluation is proposed in a concurrent work by Behnamnia et al. (2025) where a regret bound analysis of this estimator under heavy-tailed weighted reward assumption is proposed. More details about similarities and differences with this work are provided in Appendix B.
Although rich experiments are given by Li et al. (2021) for the TERM algorithm in different applications, the generalization error of the TERM has not yet been addressed for unbounded loss functions with bounded $(1+\epsilon)$-th moment for some $\epsilon\in(0,1]$.
Generalization error under unbounded loss functions: Several studies have investigated the generalization error of linear empirical risk under unbounded loss functions. Some works studied the generalization error under unbounded loss functions via the PAC-Bayesian approach. Losses with heavier tails are studied by Alquier & Guedj (2018) where probability bounds are developed. Using a different estimator than empirical risk, PAC-Bayes bounds for losses with bounded second and third moments are developed by Holland (2019). Notably, their bounds include a term that can increase with the number of samples $n$. Kuzborskij & Szepesvári (2019) and Haddouche & Guedj (2022) also provide bounds for losses with a bounded second moment. The bounds in Haddouche & Guedj (2022) rely on a parameter that must be selected before the training data is drawn. Information-theoretic bounds based on the second moment of $\sup_{h\in\mathcal{H}}|\ell(h,Z)-\mathbb{E}[\ell(h,\tilde{Z})]|$ are derived in Lugosi & Neu (2022, 2023). In contrast, our second moment assumption is more relaxed, being based on the expected version with respect to the distribution over the hypothesis set and the data-generating distribution. Furthermore, an upper bound on generalization error via VC-dimension and growth function under bounded $(1+\epsilon)$-th moment for $\epsilon\in(0,1]$ is proposed in (Cortes et al., 2019, Corollary 12) which is motivated by relative deviation generalization bounds in binary classification. In addition, the final convergence rate for unbounded loss is $O(\log(n)n^{\frac{-\epsilon}{1+\epsilon}})$ based on (Cortes et al., 2019, Corollary 12). In contrast, we derive the results for a multi-classification scenario with a convergence rate of $O(n^{\frac{-\epsilon}{1+\epsilon}})$. Our work focuses on tilted empirical risk, which is overlooked in the literature on unbounded loss functions.
8 Conclusion
@@
In this paper, we study the tilted empirical risk minimization, as proposed by Li et al. (2021). In particular, we established an upper and lower bound on the tilted generalization error of the tilted empirical risk through uniform and information-theoretic approaches, obtaining theoretical guarantees that the convergence rate is $O(n^{-\epsilon/(1+\epsilon)})$ under unbounded loss functions for negative tilt provided that $(1+\epsilon)$-th moment of loss function is bounded for some $\epsilon\in(0,1]$. We also study the tilted generalization error under distribution shift in the training dataset due to noise or outliers, where we discussed the generalization and robustness trade-off. We also explore the KL-regularized tilted empirical risk minimization, where the solution involves the tilted Gibbs posterior, and we derive a parametric upper bound on this minimization with a convergence rate of $O(n^{-\epsilon})$ under some conditions and unbounded loss function provided that $(1+\epsilon)$-th moment of loss functions is bounded for some $\epsilon\in(0,1]$.
9 Future Works
Building on our current results, we highlight several promising directions for future work:
Overparameterization Regime: An interesting avenue is to explore generalization bounds for tilted empirical risk in overparameterized models, particularly in the mean-field regime. Techniques from recent works, such as Aminian et al. (2023), may offer valuable insights for extending our results in this setting.
Removing Sample Size Assumptions for Unbounded Losses: For unbounded loss functions, we currently require a lower bound on the number of training samples to establish generalization guarantees. While we relaxed this assumption for the bounded loss case (see Appendix H), extending the analysis to the unbounded case without any sample size constraints would be a valuable extension.
Positive Tilt ($\gamma>0$): Our analysis focuses on the case of negative tilting ($\gamma<0$), assuming that the unbounded loss function has a bounded $(1+\epsilon)$-th moment for some $\epsilon\in(0,1]$. Extending the generalization analysis to positive tilting scenarios with unbounded losses remains an open and important direction.
Alternative Generalization Frameworks: Our current bounds rely on uniform and information-theoretic frameworks. However, other frameworks—such as PAC-Bayesian analysis (Catoni, 2003), algorithmic stability (Bousquet & Elisseeff, 2002b), and Rademacher complexity (Bartlett & Mendelson, 2002)—offer complementary perspectives on generalization. We derived results using these frameworks under the bounded loss function assumption in Appendix H. Investigating the applicability of these methods to tilted empirical risk under an unbounded loss function could yield deeper theoretical insights.
Acknowledgment
@@
Gholamali Aminian, Gesine Reinert and Samuel N. Cohen acknowledge the support of the UKRI Prosperity Partnership Scheme (FAIR) under EPSRC Grant EP/V056883/1 and the Alan Turing Institute. Gesine Reinert is also supported in part by EPSRC grants EP/W037211/1 and EP/R018472/1. Amir R. Asadi is supported by Leverhulme Trust grant ECF-2023-189 and Isaac Newton Trust grant 23.08(b). The authors thank the anonymous referees and the area chair for their insightful and constructive comments, which have led to an improved version of the paper. For the purpose of Open Access, the authors note that a CC BY public copyright license applies to any Author Accepted Manuscript version arising from this submission.
Impact Statement
The goal of this work is to improve our understanding of tilted empirical risk (TER) and to enhance its practical applicability by introducing a data-driven approach for hyperparameter tuning. However, we acknowledge that our current theoretical analysis, which focuses on the negative tilt regime for understanding the robustness of TER, does not capture all aspects of TER—particularly its behavior under positive tilt for biases and imbalance scenarios.
Appendix
[…]
Appendix A Summary of Notations
All notations are summarized in Table 3.
@@ Appendix A Summary of Notations
$\overline{\mathrm{R}}_{\gamma}(h,\mu^{\otimes n})$ Tilted population risk $\mathfrak{E}_{\gamma}(\mu)$ Excess risk
@@ Appendix B Other Related Works
This section details related works about tilted empirical risk minimization, the Rademacher complexity and stability and PAC-Bayesian bounds. We also compare our work with Behnamnia et al. (2025) in more details.
Generalization Error Analysis: Different approaches have been applied to study the generalization error of general learning problems under empirical risk minimization, including VC dimension-based, Rademacher complexity, PAC-Bayesian, stability and information-theoretic bounds.
Uniform Bounds: Uniform bounds (or VC bounds) are proposed by  Vapnik & Chervonenkis (1971); Bartlett et al. (1998, 2019). For any class of functions $\mathcal{F}$ of VC dimension $d$, with probability at least $1-\delta$ the generalization error is $O\big{(}(d+\log(1/\delta))^{1/2}n^{-1/2}\big{)}$. This bound depends solely on the VC dimension of the function class and on the sample size; in particular, it is independent of the learning algorithm.
Information-theoretic bounds: Russo & Zou (2019); Xu & Raginsky (2017) propose using the mutual information between the input training set and the output hypothesis to upper bound the expected generalization error. Multiple approaches have been proposed to tighten the mutual information-based bound: Bu et al. (2020) provide tighter bounds by considering the individual sample mutual information; Asadi et al. (2018); Asadi & Abbe (2020) propose using chaining mutual information; and Steinke & Zakynthinou (2020); Hafez-Kolahi et al. (2020); Aminian et al. (2020, 2021b) provide different upper bounds on the expected generalization error based on the linear empirical risk framework.
@@
Rademacher Complexity Bounds: This approach is a data-dependent method to provide an upper bound on the generalization error based on the Rademacher complexity of the function class $\mathcal{H}$, see (Bartlett & Mendelson, 2002; Golowich et al., 2018). Bounding the Rademacher complexity involves the model parameters. Typically, in Rademacher complexity analysis, a symmetrization technique is used which can be applied to the empirical risk, but not directly to the TER.
@@
Stability Bounds: Stability-based bounds for generalization error are given in Bousquet & Elisseeff (2002a); Bousquet et al. (2020); Mou et al. (2017); Chen et al. (2018); Aminian et al. (2023). For stability analysis, the key tool is Lemma 7 in Bousquet & Elisseeff (2002b), which is based on ERM linearity. Therefore, we can not apply stability analysis to TER directly.
@@
PAC-Bayesian bounds: First proposed by Shawe-Taylor & Williamson (1997); McAllester (1999) and McAllester (2003), PAC-Bayesian analysis provides high probability bounds on the generalization error in terms of the KL divergence between the data-dependent posterior induced by the learning algorithm and a data-free prior that can be chosen arbitrarily (Alquier, 2021). There are multiple ways to generalize the standard PAC-Bayesian bounds, including using information measures other than KL divergence (Alquier & Guedj, 2018; Bégin et al., 2016; Hellström & Durisi, 2020; Aminian et al., 2021b) and considering data-dependent priors (Rivasplata et al., 2020; Catoni, 2007; Dziugaite & Roy, 2018; Ambroladze et al., 2007). However, this method has not been applied to TER to provide generalization error bounds.
The aforementioned approaches are applied to study the generalization error in the linear empirical risk framework. To our knowledge, the generalization error of the tilted empirical risk minimization from information-theoretical and uniform approach perspectives has not been explored.
Comparison with Behnamnia et al. (2025): In a concurrent work in reinforcement learning, Behnamnia et al. (2025) introduce an estimator for off-policy evaluation and learning based on the log-sum-exponential (LSE) operator applied to weighted rewards. This estimator can be viewed as a negatively tilted risk function. Their focus lies in bounding the regret, defined as the gap between the value of the optimal policy and that induced by the estimator, and in deriving concentration inequalities under heavy-tailed reward distributions. We compare our work with theirs from several perspectives:
Objective Functions: Our study targets supervised learning, where the goal is to minimize a risk function such as empirical or tilted risk. In contrast, Behnamnia et al. (2025) addresses off-policy learning, aiming to maximize expected reward for a given logged data with respect to a policy, conditional distribution over actions for a given context. Furthermore, in a supervised learning scenario, the loss function is given. In contrast, in off-policy learning, the reward function is not known, and some samples of reward are given. Therefore, the learning paradigms and optimization targets differ. For more details, comparison of these scenarios see (Swaminathan & Joachims, 2015, Table 1).
Assumptions: We assume a bounded hypothesis class and heavy-tailed loss functions. In contrast, (Behnamnia et al., 2025, Assumption 5.1) considers a bounded policy class and assume heavy-tailed weighted rewards. These differences reflect the distinct nature of these tasks: off-policy and supervised learning scenarios. Note that the definition of heavy-tailed random variable is general in statistical learning theory (Resnick, 2007) and can be applied in different contexts, e.g., bandit, off-policy learning, and supervised learning.
Methodology: Despite differing goals, both works employ some common statistical tools, such as concentration inequalities. We build on analytical techniques similar to theirs, including lemmas concerning variance bounds under heavy-tailed distributions. However, our uniform bounds operate under different assumptions and are adapted to the supervised learning setting. In particular, we used Lemma C.2 and Lemma C.3– where a different proof of Lemma C.3 is proposed in (Behnamnia et al., 2025, Theorem 5.3)– to derive our upper bounds on absolute value of generalization error (Theorem 3.5) using uniform approach under heavy-tailed assumption. However, as discussed in Section 3.1, the uniform approach has inherent limitations, e.g., assuming bounded hypothesis sets which is a complexity measure in the supervised learning scenario. The limitations also apply to results in (Behnamnia et al., 2025, Theorem 5.3) with finite policy set as a complexity measure in off-policy learning. To overcome the limitations of finiteness of sets assumptions, we further develop new bounds using information-theoretic techniques.
Bounded Scenario: Both works note limitations on sample complexity under heavy-tailed assumptions. We show in Appendix H that such constraints can be lifted by assuming bounded loss functions—leading to generalization guarantees for all number of training samples.
Robustness: A key contribution of our work is the analysis of tilted empirical risk under distribution shift, a direction not explored in earlier literature to our knowledge. Our results highlight the distinct value of TER in handling training-test distribution discrepancies.
Regularization: We also explored the KL-regularized problem for tilted empirical risk, where a tighter convergence rate of $O(n^{-\epsilon})$ for tilted empirical risk was derived under a heavy-tailed assumption.
@@ Appendix C Technical Tools
We first define the functional linear derivative as in (Cardaliaguet et al., 2019).
@@
Lemma C.2 (Lemma 28 in (Lugosi & Neu, 2023)).
For $x>0$, the following inequality holds for $\epsilon\in(0,1]$,
$\exp(-x)\leq 1-x+x^{(1+\epsilon)}.$ (29)
The following lemma also appears in the concurrent work by Behnamnia et al. (2025). For the sake of completeness, we provide an alternative proof below.
@@
Suppose that $X>0$ and $\gamma<0$, then we have for $\epsilon\in(0,1]$,
$\mathrm{Var}(\exp(\gamma X))\leq 2^{\epsilon}|\gamma|^{1+\epsilon}\mathbb{E}[|X|^{1+\epsilon}].$ (30)
@@
From the variance definition, we have,
$\mathrm{Var}(\exp(\gamma X))$ $=\mathbb{E}\Big{[}\Big{(}\exp(\gamma X)-\mathbb{E}[\exp(\gamma X)]\Big{)}^{2}\Big{]}$ (31)
=$\overset{(a)}{=}\frac{1}{2}\mathbb{E}_{X,X^{\prime}}\Big{[}\Big{(}\exp(\gamma X)-\exp(\gamma X^{\prime})\Big{)}^{2}\Big{]}$ (32)
$\overset{(b)}{\leq}\frac{1}{2}\mathbb{E}_{X,X^{\prime}}\Big{[}\Big{(}\exp(\gamma X)-\exp(\gamma X^{\prime})\Big{)}^{1+\epsilon}\Big{]}$ (33)
$\overset{(c)}{\leq}\frac{1}{2}|\gamma|^{1+\epsilon}\mathbb{E}_{X,X^{\prime}}[|X-X^{\prime}|^{1+\epsilon}]$ (34)
$\overset{(d)}{\leq}\frac{1}{2}|\gamma|^{1+\epsilon}\mathbb{E}_{X,X^{\prime}}[|X+X^{\prime}|^{1+\epsilon}]$ (35)
$\overset{(e)}{\leq}2^{\epsilon}|\gamma|^{1+\epsilon}\mathbb{E}_{X}[|X|^{1+\epsilon}],$ (36)
where (a) follows from variance representation for i.i.d copy $X$ and $X^{\prime}$, (b) follows from the fact that $|\exp(\gamma X)-\exp(\gamma X^{\prime})|\leq 1$, (c) follows from the fact that $\exp(\gamma x)$ is Lipschitz for $\gamma<0$, with parameter $|\gamma|$, therefore, we have $|\exp(\gamma x)-\exp(\gamma x^{\prime})|\leq|\gamma||x-x^{\prime}|$, (d) follows the positiveness of $X$ and $X^{\prime}$ and (e) follows from Jensen-inequality. ∎
@@
For a positive random variable, $Z>0$ and $\epsilon\in(0,1]$, suppose $\mathbb{E}[Z^{1+\epsilon}]\leq\eta$ Then, the following inequality holds,
$\mathbb{E}[Z]\leq\eta^{1/{1+\epsilon}}.$
@@
Suppose $\mathbb{E}[|X|^{1+\epsilon}]<\infty$. Then, for $X>0$ and $\gamma<0$ the following inequality holds,
$0\leq\mathbb{E}[X]-\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}\leq|\gamma|^{\epsilon}\mathbb{E}[X^{1+\epsilon}].$
@@
The left inequality follows from Jensen’s inequality applied to $f(x)=\log{(x)}$. For the right inequality, from Lemma C.2 for $\epsilon\in(0,1]$
$e^{\gamma X}\leq 1+\gamma X+|\gamma|^{1+\epsilon}X^{1+\epsilon}.$ (39)
Therefore, we have,
$\begin{split}\frac{1}{\gamma}\log{\mathbb{E}[e^{\gamma X}]}&\geq\frac{1}{\gamma}\log{\mathbb{E}\left[1+\gamma X+|\gamma|^{1+\epsilon}|X|^{1+\epsilon}\right]}\\ &=\frac{1}{\gamma}\log{\left(1+\gamma\mathbb{E}[X]+|\gamma|^{1+\epsilon}\mathbb{E}[|X|^{1+\epsilon}]\right)}\\ &\geq\frac{1}{\gamma}\left(\gamma\mathbb{E}[X]+|\gamma|^{1+\epsilon}\mathbb{E}[|X|^{1+\epsilon}]\right)\\ &=\mathbb{E}[X]-|\gamma|^{\epsilon}\mathbb{E}[X^{1+\epsilon}].\end{split}$
@@
Lemma C.12 (Talagrand’s contraction lemma (Shalev-Shwartz & Ben-David, 2014)).
@@
Lemma C.14 (Lemma 1 in (Xu & Raginsky, 2017)).
@@
Lemma C.16 (Kantorovich-Rubenstein duality of total variation distance, see  (Polyanskiy & Wu, 2022)).